A rectangular table top has a perimiter of 18 inches and an area of 20 square inches. FInd its dimensions
Answer by nerdybill(7384)

A rectangular table top has a perimiter of 18 inches and an area of 20 square inches. FInd its dimensions
.
Let W = width
and L = length
.
from perimeter:
2(W+L)=18
W+L = 9 (equation 1)
.
from area:
WL = 20 (equation 2)
.
solve equation 1 for L:
W+L = 9
L = 9-W
substitute into equation 2 and solve for W:
WL = 20
W(9-W) = 20
9W-W^2 = 20
9W = W^2+20
0 = W^2-9W+20
0 = (W-4)(W-5)
W = {4,5}
.
Dimensions are: 4 inches by 5 inches